OpenStudy (anonymous):

3/4(8k-4)=-3+6k Identity or no solution? :)

OpenStudy (anonymous):

is this yor question \[\frac{ 3 }{ 4 }(8k-4)=-3+6k\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

were ya sleeping lol

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

\[3*2k-1=-3+6k\] 6k-1= -3 +6k so it has got no solution , don sleep cat lolhahaha

OpenStudy (anonymous):

Im just checking my answer if im right

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

Your answer is not same as mine

OpenStudy (anonymous):

cus ya solved wrong

OpenStudy (anonymous):

Hmm, i'm not sure who's right :) how did you get the answer for 3/4(8k-4)?

hero (hero):

\[\frac{ 3 }{ 4 }(8k-4)=-3+6k \\\frac{ 3 \dot\ 4 }{ 4 }(2k-1)=-3+6k \\3(2k - 1) = -3 + 6k \\6k - 3 = -3 + 6k \\6k - 3 = 6k - 3 \] Identity

OpenStudy (anonymous):

Oh thanks, we have the same answer :) i'm just confused for the answer

OpenStudy (anonymous):

ma honest mistake

hero (hero):

6k - 3 = 6k - 3 6k - 6k = 3 - 3 0 = 0 Identity